# SAGe: A Lightweight Algorithm-Architecture Co-Design for Alleviating Data Preparation Overheads in Large-Scale Genome Analysis

### v1

@@ 2. Background
2.1. SSD Organization
Fig. 1 shows a modern NAND flash-based SSD’s organization.
Figure 1. Organizational overview of a modern SSD.
1
NAND Flash Memory. One or more NAND packages share a channel to communicate with the SSD controller. Each NAND package contains one or more (typically up to 4) flash dies or chips that share the NAND package’s I/O pins.
2
SSD Controller. The SSD controller has two key components. First, multiple embedded cores execute the flash translation layer (FTL). Second, per-channel flash controllers perform request handling and error correction for reliability.
3
DRAM. SSDs utilize an internal DRAM to improve SSD performance. It uses the DRAM to store (i) the logical-to-physical (L2P) mappings, (ii) recently-accessed data, and (iii) metadata for management tasks.
@@ 2.2. Genomic Workflow
Given a genomic sample, a typical workflow consists of three key steps (alser2020technology, ; alser2020accelerating, ; berger2023navigating, ): sequencing, basecalling, and analysis. The first step, sequencing, involves reading the DNA from a biological sample as digital signals. Due to limitations in current sequencing technologies, many organisms’ entire genomes cannot be read as complete, error-free sequences. Instead, these technologies produce many shorter sequences called reads that are randomly sampled from different locations in the genome, where the average number of reads per location is called the sequencing depth. Deeper sequencing provides more information per location, allowing later analysis steps to better distinguish between biological signals and sequencing errors. State-of-the-art sequencing technologies are distinguished by the characteristics of the reads they produce. Examples include Illumina sequencers that produce short (corresponding to 150 characters) reads with high accuracy ($\sim$99.9%) (glenn2011field, ; goodwin2016coming, ; quail2012tale, ; kchouk2017generations, ; pfeiffer2018systematic, ), PacBio HiFi sequencing (wenger2019accurate, ; hon2020highly, ; amarasinghe2020opportunities, ) and recent ONT sequencers (sereika2022oxford, ; ni2023benchmarking, ) that produce long (1k to 25k) reads with intermediate accuracy ($\sim$99%) (hon2020highly, ; ni2023benchmarking, ), and older ONT sequencers that produce long reads (1.5k to 7.5k (ni2023benchmarking, )) with lower accuracy ($\sim$85-94%) (wang2021nanopore, ).
@@
The second step, basecalling, converts the raw digital signals from a sequencer into strings. The signal from each DNA read is converted into three strings (cock2009sanger, ): (i) a header containing identifiers for the sequencing experiment and the specific read, (ii) the DNA encoded using an alphabet representing the nucleotides (also called base pairs) A, C, G, T (alongside N to represent an uncertain position), and (iii) a quality score (ewing1998base, ) for each nucleotide, encoding its probability of being incorrect using ASCII characters. The strings from all reads in a sample are then written sequentially [^2] This format is called FASTQ (cock2009sanger, ), the most commonly used read set format (illuminafastq, ). and stored in a compressed form (§2.3) for future analysis steps.
@@
The third step is genome analysis on a collection of read sets. Genomic read sets can be obtained by, for example, (i) sequencing the DNA in a sample and storing the reads into the SSD of a sequencing machine (gamaarachchi2022fast, ; minion21, ) or (ii) downloading the read sets to a local SSD from a repository (leinonen2010sequence, ; bokulich2020measuring, ). A typical analysis workflow quantifies significant sequence mismatches between the reads and a reference genome. This typically involves two key tasks: First, the computationally-expensive read mapping process finds the potential matching locations of reads in a reference genome. Second, variant calling uses the mapping information to determine which observed mismatches are likely to be true biological sequence variation (and not sequencing errors). [^3] Analysis workflows access all DNA bases from the reads, but only a small fraction of the corresponding quality scores. This is because read mapping, while using all bases, typically ignores quality scores (li2018minimap2, ; vasimuddin2019efficient, ). The subsequent variant calling step only needs quality scores from the positions surrounding mismatches (Poplin2018, ; li2008mapping, ).
@@
In many cases, the analysis step for a single read set can be performed many times and at different times. Some applications (e.g., measuring the genetic diversity in a population (kostlbacher2021pangenomics, ; vandorp2020emergence, )) require analyzing the genetic differences between a read set belonging to an individual and many reference genomes. Other applications require repeating the analysis step many times (e.g., with different read mapping parameters (khayat2021hidden, ; berger2023navigating, ) or updated reference genomes (chen2024improved, ; aganezov2022complete, ; berger2023navigating, ; Siren2024, )) to improve accuracy. Due to the criticality of the analysis step and the challenges of analyzing large amounts of genomic data on conventional systems, a large body of works (e.g., (mutlu2023accelerating, ; alser2022molecules, ; lou2020helix, ; lou2018brawl, ; shahroodi2023swordfish, ; saavedra2020mining, ; markus2020benchmarking, ; subramaniyan2021accelerated, ; huangfu2018radar, ; khatamifard2021genvom, ; cali2020genasm, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; turakhia2018darwin, ; fujiki2018genax, ; madhavan2014race, ; cheng2018bitmapper2, ; houtgast2018hardware, ; houtgast2017efficient, ; zeni2020logan, ; ahmed2019gasal2, ; nishimura2017accelerating, ; de2016cudalign, ; liu2015gswabe, ; liu2013cudasw++, ; liu2009cudasw++, ; liu2010cudasw++, ; wilton2015arioc, ; goyal2017ultra, ; chen2016spark, ; chen2014accelerating, ; chen2021high, ; fujiki2020seedex, ; banerjee2018asap, ; fei2018fpgasw, ; waidyasooriya2015hardware, ; chen2015novel, ; rucci2018swifold, ; haghi2021fpga, ; li2021pipebsw, ; ham2020genesis, ; ham2021accelerating, ; wu2019fpga, )) accelerate genome analysis tasks.
@@ 2.3. Storing and Accessing Genomic Data
Driven by continuing exponential drops in DNA sequencing costs (nhgri, ; pedro2021integration, ), genomic sequence data volumes are growing by an order of magnitude every few years, approaching and expected to exceed the data volumes generated by various social media platforms (stephens2015big, ; srastats, ; katz2021sra, ; enastats, ). Due to the importance of efficiently storing large volumes of genomic data in a space-efficient manner, there exist many compression techniques (e.g., (chandak2018spring, ; chandak2017compression, ; kokot2022colord, ; Meng2023, ; roguski2018fastore, ; dufort2021renano, ; kowalski2019pgrc, ; dufort2020enano, )) specialized for genomic data to achieve higher compression ratios (e.g., typically ranging from $\sim$2 to $\sim$40 (chandak2018spring, ; roguski2018fastore, ; dufort2020enano, ; dufort2021renano, )) than general-purpose compression techniques (e.g., typically ranging from $\sim$2 to $\sim$(chandak2018spring, )). These genomics-specific compression techniques leverage the typically high degree of data redundancy in genomic datasets (especially at high sequencing depths, as explained in §2.2) to maximize compression ratios.
Fig. 2 shows an example of a typical genomics-specific compression technique. To store each read set in a compressed way, genomics-specific compression tools typically represent the read set with a 1 consensus sequence and 2 the mismatches of each read in the read set compared to that consensus (chandak2017compression, ; dufort2021renano, ; kokot2022colord, ; kowalski2019pgrc, ). A consensus sequence is an approximation of the sequenced organism’s genome, and can either be a user-provided reference genome (dufort2021renano, ) or a de-duplicated string derived from the reads, representing the most likely character at each location (chandak2017compression, ; kowalski2019pgrc, ). [^4] Note that it is not sufficient to store only the consensus sequence to represent the read set because the consensus sequence alone does not capture the full variability in the original reads. Individual reads often contain unique differences, including sequencing errors (goodwin2016coming, ; wang2021nanopore, ) or biological variation (e.g., when multiple copies of the same chromosome from each parent are present in the DNA sample (Cheng2022, )). Given this consensus sequence, a lossless encoding of a read consists of 1 its matching position in the consensus sequence (many compression tools, e.g., (chandak2017compression, ; Meng2023, ; roguski2018fastore, ; kowalski2019pgrc, ), store matching positions in a sorted manner based on their order of appearance in the consensus sequence, which enables delta encoding to reduce storage requirements), 2 mismatch positions (delta-encoded), 3 mismatch bases and types (i.e., substitution, insertion, deletion), and 4 read length (relevant for long-read sets since they have variable read lengths). This encoding of the reads is better compressible using general-purpose compressors (chandak2017compression, ). State-of-the-art genomic compression techniques further compress the mismatch information using various general compression techniques.
@@
Note that identifying matching positions against the consensus sequence while compressing a read set does not eliminate the need for read mapping during subsequent analysis stages (Siren2024, ; Sedlazeck2018, ; Abel2020, ). This is due to three reasons. First, when compressing and storing a sequenced and basecalled read set, the exact reference genomes required for potential future read mapping analyses may not be known in advance (Siren2024, ). Second, a read set may undergo read mapping multiple times or at different times (as explained in §2.2). Therefore, simply having the matches against a single consensus sequence during compression is insufficient. Third, during compression, only one matching position for each read is stored (to minimize data size), whereas various read mapping use cases require many matching positions of each read (Sedlazeck2018, ; Abel2020, ; li2018minimap2, ).
@@ 3. Motivation
We perform experimental studies to understand the impact of preparing genomic data (i.e., accessing the data from storage and decompressing and reformatting it) for a genome analysis accelerator on end-to-end performance.
@@ 3.1. Methodology
Genome Analysis Accelerator. We use a state-of-the-art hardware accelerator (chen2023gem, ) for read mapping. To find the read mapping throughput, we use the throughput reported by the original paper.
@@
To find the performance of data access from the SSD and decompression, we use a state-of-the-art SSD (samsungPM1735, ) with PCIe Gen4 (PCIE4, ) and a high-end server with AMD EPYC 7742 CPU (amdepyc, ) with 128 physical cores and 1.5-TB DDR4 DRAM. We use the best-performing thread count (i.e., the thread count after which adding more threads does not improve performance) for Spring (chandak2018spring, ) and pigz (adler2015pigz, ).
End-to-End Performance. We find the end-to-end throughput of the genome analysis task based on the throughput of preparing data for the accelerator and the throughput of the read mapping accelerator itself. I/O operations, decompression and reformatting, and read mapping operate in a pipelined manner and in batches (i.e., when decompressing and reformating reads in $batch\#i$, the mapper analyzes $batch\#i-1$), which enables partially overlapping parts of these steps. Further methodology details are in §6.
Evaluated Configurations. To show the impact of data preparation, we analyze configurations with different initial states for the input read set, each incurring different decompression and I/O overheads: (i) Cmprs1+I/O and (ii) Cmprs2+I/O with data stored in pigz- and Spring-compressed formats, respectively, within the storage system. (iii) Cmprs1+NoI/O and (iv) Cmprs2+NoI/O with data stored in pigz- and Spring-compressed formats, respectively, but in an idealized scenario with zero performance overhead due to storage I/O; (v) NoCmprs+I/O with data stored in an uncompressed and desired format for the accelerator in the storage system; and (vi) NoCmprs+NoI/O with data stored in an uncompressed format and zero performance overhead caused by storage I/O. This configuration assumes that the data is decompressed, in the desired format, and ready to be fed to the accelerator. We show each system’s throughput normalized to the throughput of NoCmprs+NoI/O.
@@ 3.2. Observations
Figure 3. Read mapping throughput with six configurations for a read set’s initial state.
@@
Fig. 3 shows end-to-end throughput of preparing data and performing read mapping on it. We make four key observations. First, the overhead of preparing data for the accelerator significantly hurts end-to-end performance. Cmprs1+I/O and Cmprs2+I/O, which suffer from throughput loss due to I/O and decompression overheads, lead to 51.5$\times$ and 27.0$\times$ slowdown compared to NoCmprs+NoI/O, respectively. Second, the overhead of decompressing data to the desired format is larger than the I/O overhead. Removing the I/O overhead (in the Cmprs1+NoI/O and Cmprs2+NoI/O configurations) does not improve performance compared to Cmprs1+I/O and Cmprs2+I/O, as the decompression step dominates the end-to-end throughput of the data preparation pipeline. Third, decompression overhead has a larger impact on end-to-end throughput than the I/O overhead of reading larger uncompressed data from the storage system. Although NoCmprs+I/O requires reading larger uncompressed data from storage and thus incurs larger I/O overhead than Cmprs1+I/O and Cmprs2+I/O, it still achieves higher throughput than them (by 20.3$\times$ and 10.6$\times$, respectively). However, it is inefficient and unsustainable to store large volumes of genomic data uncompressed in the storage system. Fourth, I/O overhead, although less severe than the overhead of decompressing data into the desired format, still considerably affects end-to-end performance. NoCmprs+I/O leads to 2.5$\times$ slowdown compared to NoCmprs+NoI/O. Note that this I/O overhead is independent of the relative size of the read set and main memory since this I/O overhead is due to the need to move large amounts of low reuse data from the storage system to the read mapping accelerator (mansouri2022genstore, ). As shown in Fig. 3, the I/O overhead exists even when the main memory is larger than the read set.
3.3. Our Goal
@@
Based on our observations, we conclude that data preparation leads to large overheads in the end-to-end performance of genome analysis, and can significantly diminish the potential benefits of genome analysis accelerators. Therefore, to be able to actually harness the full benefits of the extensive efforts on accelerating genome analysis, there is a critical need for a design that effectively addresses data preparation overheads while meeting three key requirements: (ihigh performance and energy efficiency for data preparation, (iihigh compression ratios, comparable to state-of-the-art genomic compression algorithms, and (iii) be lightweight to enable seamless integration with a broad range of genome analysis systems, regardless of where they are implemented.
@@
Unfortunately, it is challenging to optimize both the performance of preparing data and the space-efficiency of storing data while maintaining low hardware costs. This is particularly challenging when targeting integration with genome analysis systems in resource-constrained environments. For example, while near data processing (NDP) leads to significant benefits in genome analysis (wu2021sieve, ; shahroodi2022krakenonmem, ; shahroodi2022demeter, ; dashcam23micro, ; hanhan2022edam, ; zou2022biohd, ; cali2020genasm, ; huangfu2018radar, ; khatamifard2021genvom, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; Zhang_2023_alignerD, ; mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, ) by fundamentally alleviating data movement overheads when analyzing large-scale genomic data, such designs are often implemented in resource-constrained environments (e.g., within the memory system or storage device), which makes expensive decompression techniques prohibitive to deploy. Consider NDP accelerators implemented inside the storage system (e.g., (mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, )), which provide large performance and energy benefits for genome analysis by (i) fundamentally alleviating data-movement overhead from the storage system (and thus help alleviate the I/O overhead in data preparation), and (ii) reducing the application’s burden from the rest of the system (i.e., computational units and main memory). Without the capability to prepare data for analysis inside the SSD’s resource-constrained environment, all such works would either necessitate data to be stored in an uncompressed way, which is inefficient, or require external decompression, which completely undermines the fundamental benefits of NDP.
@@
Some prior works (e.g., (guo2013gpu, ; zhao2017streaming, ; jiang2021exma, ; arram2015fpga, ; chen2023efficient, ; leavline2013hardware, ; qiao2019fpga, ; wang2018accelerating, )) accelerate certain computational kernels, such as BWT (qiao2019fpga, ; guo2013gpu, ; zhao2017streaming, ), FM-index search (jiang2021exma, ; arram2015fpga, ), or LZMA (chen2023efficient, ; leavline2013hardware, ), which are widely used in many state-of-the-art genomics-specific compression tools (e.g., (lan2021genozip, ; Meng2023, ; kowalski2019pgrc, ; chandak2017compression, )). Despite their benefits, these works have two limitations. First, existing techniques mainly rely on many random accesses to large data structures (due to random look-ups for matching patterns in these structures), which makes their adoption inefficient in resource-constrained environments. For example, some works (e.g., (zhao2017streaming, ; wang2018accelerating, ; dragenora, )) demand high DRAM bandwidth. Some works require expensive hardware components, such as large buffers (qiao2019fpga, ), DRAM capacity (jiang2021exma, ), or large computation resources (arram2015fpga, ; guo2013gpu, ; leavline2013hardware, ). Therefore, these techniques are prohibitive to support in resource-constrained environments, such as portable genome analysis devices or NDP accelerators. [^6] For example, an NDP accelerator inside the storage system is implemented inside the resource-constrained environment of an SSD, with the limited-bandwidth single-channel internal DRAM (zou2022assasin, ) with small capacity (e.f., 1-GB DRAM for each TB capacity. For example, a 4-GB LPDDR4 DRAM is used in a modern 4-TB SSD (samsung860pro, ). While one can reduce the required buffer/DRAM spaces by decreasing window sizes, such an approach consequently lowers compression ratios (qiao2019fpga, ), which is not acceptable for adoption for genomic data. Second, in addition to their adoption limitations, these tools do not fully alleviate the end-to-end overhead of data preparation (as shown in §7) since they do not consider all of the steps involved in data preparation.
Our goal in this work is to improve the performance and energy efficiency of genome analysis by mitigating data preparation overheads, while maintaining compression ratios comparable to state-of-the-art genomics-specific compressors, and ensuring a lightweight design that can be seamlessly integrated with a broad range of genome analysis systems, such as general-purpose systems, specialized accelerators (mutlu2023accelerating, ; alser2022molecules, ; lou2020helix, ; lou2018brawl, ; shahroodi2023swordfish, ; saavedra2020mining, ; markus2020benchmarking, ; subramaniyan2021accelerated, ; huangfu2018radar, ; khatamifard2021genvom, ; cali2020genasm, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; turakhia2018darwin, ; fujiki2018genax, ; madhavan2014race, ; cheng2018bitmapper2, ; houtgast2018hardware, ; houtgast2017efficient, ; zeni2020logan, ; ahmed2019gasal2, ; nishimura2017accelerating, ; de2016cudalign, ; liu2015gswabe, ; liu2013cudasw++, ; liu2009cudasw++, ; liu2010cudasw++, ; wilton2015arioc, ; goyal2017ultra, ; chen2016spark, ; chen2014accelerating, ; chen2021high, ; fujiki2020seedex, ; banerjee2018asap, ; fei2018fpgasw, ; waidyasooriya2015hardware, ; chen2015novel, ; rucci2018swifold, ; haghi2021fpga, ; li2021pipebsw, ; ham2020genesis, ; ham2021accelerating, ; wu2019fpga, ), and near-data processing architectures (e.g.,in main memory (wu2021sieve, ; shahroodi2022krakenonmem, ; shahroodi2022demeter, ; dashcam23micro, ; hanhan2022edam, ; zou2022biohd, ; cali2020genasm, ; huangfu2018radar, ; khatamifard2021genvom, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; Zhang_2023_alignerD, ) or storage system (mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, )).

### v2

@@ 2.1. Genomic Workflow
Given a genomic sample, a typical workflow consists of three key steps (alser2020technology, ; alser2020accelerating, ; berger2023navigating, ): sequencing, basecalling, and analysis. The first step, sequencing, involves reading the DNA from a biological sample as digital signals. Due to limitations in current sequencing technologies, many organisms’ entire genomes cannot be read as complete, error-free sequences. Instead, these technologies produce many shorter sequences called reads that are randomly sampled from different locations in the genome. The average number of reads per location is called the sequencing depth. Deeper sequencing provides more information per location, allowing later analysis steps to better distinguish between biological signals and sequencing errors. Sequencers are distinguished by the characteristics of their reads. Examples include sequencers that produce short (75 to 300 characters) reads with high accuracy ($\sim$99.9%) (glenn2011field, ; goodwin2016coming, ; quail2012tale, ; kchouk2017generations, ; pfeiffer2018systematic, ), long (1k to 25k) reads with intermediate accuracy ($\sim$99%) (hon2020highly, ; ni2023benchmarking, ; wenger2019accurate, ; amarasinghe2020opportunities, ; sereika2022oxford, ), and older sequencers with long reads (1.5k to 7.5k (ni2023benchmarking, )) and lower accuracy ($\sim$85-94%) (wang2021nanopore, ).
@@
The second step, basecalling, converts the raw digital signals from a sequencer into strings. The signal from each DNA read is converted into three strings (cock2009sanger, ): (i) a header containing identifiers for the sequencing experiment and the read, (ii) the DNA encoded using an alphabet representing the nucleotides (also called base pairs) A, C, G, T (alongside N to represent an uncertain position), and (iii) a quality score (ewing1998base, ) for each nucleotide, encoding its probability of being incorrect using ASCII characters. The strings from all reads in a sample are then written sequentially [^2] This format is called FASTQ (cock2009sanger, ), the most commonly used read set format (illuminafastq, ). and stored in a compressed form (§2.2) for future analysis steps.
@@
The third step is genome analysis on a collection of read sets. Read sets can be obtained by, for example, (i) sequencing the DNA in a sample and storing the reads into the SSD of a sequencer (gamaarachchi2022fast, ; minion21, ) or (ii) downloading the read sets to a local SSD from a repository (leinonen2010sequence, ; bokulich2020measuring, ). A typical analysis workflow quantifies significant mismatches between the reads and a reference genome. This typically involves two key tasks: First, the computationally-expensive read mapping process finds the potential matching locations of reads in a reference genome. Second, variant calling uses the mapping information to determine which observed mismatches are likely to be true biological variation (and not sequencing errors). [^3] Analysis workflows access all DNA bases from the reads, but only a small fraction of the corresponding quality scores. This is because read mapping, while using all bases, typically ignores quality scores (li2018minimap2, ; vasimuddin2019efficient, ). The subsequent variant calling step only needs quality scores from the positions surrounding mismatches (Poplin2018, ; li2008mapping, ).
@@
In many cases, the analysis step for a single read set can be performed many times and at different times. Some applications (e.g., measuring the genetic diversity in a population (kostlbacher2021pangenomics, ; vandorp2020emergence, )) require analyzing the differences between a read set belonging to an individual and many reference genomes. Other applications require repeating the analysis many times (e.g., with different read mapping parameters (khayat2021hidden, ; berger2023navigating, ) or updated reference genomes (chen2024improved, ; aganezov2022complete, ; berger2023navigating, ; Siren2024, )) to improve accuracy. Due to the criticality of the analysis step and the challenges of analyzing large amounts of genomic data on conventional systems, a large body of works (e.g., (mutlu2023accelerating, ; alser2022molecules, ; lou2020helix, ; lou2018brawl, ; shahroodi2023swordfish, ; saavedra2020mining, ; markus2020benchmarking, ; subramaniyan2021accelerated, ; huangfu2018radar, ; khatamifard2021genvom, ; cali2020genasm, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; turakhia2018darwin, ; fujiki2018genax, ; madhavan2014race, ; cheng2018bitmapper2, ; houtgast2018hardware, ; houtgast2017efficient, ; zeni2020logan, ; ahmed2019gasal2, ; nishimura2017accelerating, ; de2016cudalign, ; liu2015gswabe, ; liu2013cudasw++, ; liu2009cudasw++, ; liu2010cudasw++, ; wilton2015arioc, ; goyal2017ultra, ; chen2016spark, ; chen2014accelerating, ; chen2021high, ; fujiki2020seedex, ; banerjee2018asap, ; fei2018fpgasw, ; waidyasooriya2015hardware, ; chen2015novel, ; rucci2018swifold, ; haghi2021fpga, ; li2021pipebsw, ; ham2020genesis, ; ham2021accelerating, ; wu2019fpga, )) accelerate genome analysis tasks.
@@ 2.2. Storing and Accessing Genomic Data
Driven by continuing exponential drops in sequencing costs (nhgri, ; pedro2021integration, ), genomic data volumes are growing by an order of magnitude every few years, approaching and expected to exceed the data volumes generated by various social media platforms (stephens2015big, ; srastats, ; katz2021sra, ; enastats, ). Due to the importance of storing genomic data space-efficiently, there exist many compression techniques (e.g., (chandak2018spring, ; chandak2017compression, ; kokot2022colord, ; Meng2023, ; roguski2018fastore, ; dufort2021renano, ; kowalski2019pgrc, ; dufort2020enano, )) specialized for genomic data, achieving higher compression ratios (e.g., typically from $\sim$2 to $\sim$40 (chandak2018spring, ; roguski2018fastore, ; dufort2020enano, ; dufort2021renano, )) than general-purpose compression techniques (e.g., typically from $\sim$2 to $\sim$(chandak2018spring, )). These genomic compressors leverage the typically high degree of redundancy in genomic datasets (especially at high sequencing depth).
@@
Fig. 2 shows an example of a typical genomics-specific compression technique. Genomic compressors typically represent each read set with a 1 consensus sequence and 2 the mismatches of each read in the read set compared to that consensus (chandak2017compression, ; dufort2021renano, ; kokot2022colord, ; kowalski2019pgrc, ). A consensus sequence is an approximation of the organism’s genome, and can either be a user-provided reference genome (dufort2021renano, ) or a de-duplicated string derived from the reads, representing the most likely character at each location (chandak2017compression, ; kowalski2019pgrc, ). [^4] Note that it is not sufficient to store only the consensus sequence to represent the read set because the consensus sequence alone does not capture the full variability in the original reads. Individual reads often contain unique differences, including sequencing errors (goodwin2016coming, ; wang2021nanopore, ) or biological variation (e.g., when multiple copies of the same chromosome from each parent are present in the DNA sample (Cheng2022, )). Given this consensus, a lossless encoding of a read consists of 1 its matching position in the consensus (many compressors, e.g., (chandak2017compression, ; Meng2023, ; roguski2018fastore, ; kowalski2019pgrc, ), store matching positions based on their order of appearance in the consensus, which enables delta-encoding to reduce storage requirements), 2 mismatch positions (delta-encoded), 3 mismatch bases and types (i.e., substitution, insertion, deletion), and 4 read length (for long-reads since they have variable lengths). This encoding of the reads is then more compressible using general compressors (chandak2017compression, ). State-of-the-art genomic compressors further compress the mismatch information using various general compression techniques.
Note that identifying matching positions against the consensus during compression does not eliminate the need for read mapping during subsequent analysis stages (Siren2024, ; Sedlazeck2018, ; Abel2020, ). This is due to three reasons. First, during compression, the exact reference genomes required for potential future mapping analyses may not be known in advance (Siren2024, ). Second, a read set may undergo mapping multiple times or at different times (as explained in §2.1). Therefore, simply having the matches against a single consensus is insufficient. Third, during compression, only one matching position for each read is stored (to minimize size), whereas various mapping use cases require many matching positions of each read (Sedlazeck2018, ; Abel2020, ; li2018minimap2, ).
@@ 3. Motivation
We perform experimental studies to demonstrate the impact of the data preparation bottleneck on the end-to-end performance of a state-of-the-art genome analysis accelerator.
@@ 3.1. Methodology
Genome Analysis Accelerator. We use a state-of-the-art hardware accelerator (chen2023gem, ) for read mapping. We use the read mapping throughput reported by the original paper.
@@
To find the performance of accessing the compressed data and decompressing it, we use a high-end server with AMD EPYC 7742 CPU (amdepyc, ) with 128 physical cores and 1.5-TB DDR4 DRAM with eight channels, and an SSD (samsungPM1735, ) with PCIe Gen4 (PCIE4, ). We use the best-performing thread count (i.e., after which adding more threads does not improve performance) for Spring (chandak2018spring, ) and pigz (adler2015pigz, ).
End-to-End Performance. We find end-to-end throughput based on the throughput of preparing data for the accelerator and the throughput of the read mapping accelerator itself. I/O operations, decompression, and read mapping operate in a pipelined manner and in batches (i.e., when decompressing reads in $batch\#i$, the mapper analyzes $batch\#i-1$), which enables partially overlapping parts of these steps. Further methodology details are in §7.
Evaluated Configurations. To show the impact of data preparation, we analyze configurations with different initial states for the input read set, each incurring different decompression and I/O overheads: (i) Cmprs1+I/O and (ii) Cmprs2+I/O with data stored in pigz- and Spring-compressed formats, respectively, within the storage system. (iii) Cmprs1+NoI/O and (iv) Cmprs2+NoI/O with data stored in pigz- and Spring-compressed formats, respectively, but in an idealized scenario with zero performance overhead due to storage I/O; (v) NoCmprs+I/O with data stored in an uncompressed and desired format for the accelerator in the storage system; and (vi) NoCmprs+NoI/O with data stored in an uncompressed format and zero performance overhead caused by storage I/O, ready to be fed to the accelerator. We show each system’s throughput normalized to the throughput of NoCmprs+NoI/O.
@@ 3.2. Observations
Figure 3. End-to-end throughput with six configurations for a read set’s initial state.
@@
Fig. 3 shows the end-to-end throughput of preparing data and performing read mapping on it. We make four key observations. First, data preparation greatly diminishes end-to-end performance. Cmprs1+I/O and Cmprs2+I/O lead to 51.5$\times$ and 27.0$\times$ slowdown compared to NoCmprs+NoI/O, respectively. Second, the overhead of decompressing data to the desired format is larger than the I/O overhead. Removing I/O overhead (in the Cmprs1+NoI/O and Cmprs2+NoI/O configurations) does not improve performance compared to Cmprs1+I/O and Cmprs2+I/O, as the decompression step dominates the end-to-end throughput of the pipeline. Third, decompression overhead has a larger impact on end-to-end throughput than the I/O overhead of reading larger uncompressed data from the storage system. Although NoCmprs+I/O requires reading larger uncompressed data from storage and thus incurs larger I/O overhead than Cmprs1+I/O and Cmprs2+I/O, it still achieves higher throughput than them (by 20.3$\times$ and 10.6$\times$, respectively). However, it is inefficient and unsustainable to store large volumes of genomic data uncompressed in the storage system. Fourth, I/O overhead, although less severe than decompression overhead, can still considerably affect end-to-end performance. NoCmprs+I/O leads to 2.5$\times$ slowdown compared to NoCmprs+NoI/O. Note that this I/O overhead is independent of the relative size of the read set and main memory, since this overhead stems from the need to move large amounts of low reuse data from the storage system to the accelerator. As shown in Fig. 3, this I/O overhead exists even when the main memory is larger than the read set. While various prior works (e.g., (mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, )) aim to alleviate this I/O overhead with near-data processing (NDP) inside the storage system, these techniques cannot work when data is stored compressed in storage.
3.3. Challenges and Goal
@@
Our observations in Figs. 1 and 3 demonstrate that as accelerators dramatically speed up genome analysis, data preparation emerges as a critical bottleneck in the execution pipeline. Therefore, there is a critical need to effectively mitigate this bottleneck while meeting three key requirements: (ihigh performance and energy efficiency, (iihigh compression ratios, comparable to state-of-the-art genomic compression algorithms, and (iii) be lightweight to enable seamless integration with a broad range of genome analysis systems (e.g., in FPGAs, GPUs, ASICs, portable devices, or NDP systems).
@@
Challenges. Optimizing both data preparation performance and achieving high compression ratios (on par with genomic compressors) at low hardware costs is challenging. This challenge is particularly exacerbated when aiming for integration with genome analysis systems in hardware resource-constrained environments. Examples include portable genomics devices (e.g., (SmidgION, ; MinIONMk1CO, ; palatnick2020igenomics, ; Ballard2018, ; Oehler2023, )), which are critical to facilitating the wide adoption of genomics, or various NDP accelerators (e.g., (wu2021sieve, ; shahroodi2022krakenonmem, ; shahroodi2022demeter, ; dashcam23micro, ; hanhan2022edam, ; zou2022biohd, ; cali2020genasm, ; huangfu2018radar, ; khatamifard2021genvom, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; Zhang_2023_alignerD, ; mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, )), which lead to large benefits by mitigating data movement overheads of analyzing large-scale genomic data, but are typically implemented in constrained settings (e.g., within the memory system or storage device).
@@
Some prior works accelerate certain computational kernels, such as BWT (qiao2019fpga, ; guo2013gpu, ; zhao2017streaming, ), FM-index search (jiang2021exma, ; arram2015fpga, ), or LZMA (chen2023efficient, ; leavline2013hardware, ), which are widely used in many state-of-the-art genomics-specific compression tools (e.g., (lan2021genozip, ; Meng2023, ; kowalski2019pgrc, ; chandak2017compression, )). Despite their benefits, these works have two limitations. First, they are unsuitable for adoption in resource-constrained environments. This is because these techniques mainly rely on many random accesses to large data structures (due to random look-ups for matching patterns in these structures), which in turn leads to high utilization of expensive resources. For example, some works (e.g., (zhao2017streaming, ; wang2018accelerating, ; dragenora, )) demand high DRAM bandwidth. Some works require expensive hardware components, such as large buffers (qiao2019fpga, ), DRAM capacity (jiang2021exma, ), or large computation resources (arram2015fpga, ; guo2013gpu, ; leavline2013hardware, ). Second, in addition to the adoption limitations, these tools do not fully mitigate the end-to-end overhead of data preparation (as shown in §8) since they do not consider all of the steps involved in data preparation.
To further underscore the challenge of seamless integration in resource-constrained environments, we showcase an example. Consider an NDP accelerator implemented inside the storage system (e.g., (mansouri2022genstore, ; abakus23taco, ; megis, ; jun2016storage, )), which provides large benefits by alleviating data movement overhead from storage and reducing the application’s burden from the rest of the system. Without the capability to prepare data inside the SSD, all such works would either necessitate data to be stored uncompressed, which is inefficient, or require decompression outside the SSD, which completely undermines the fundamental benefits of NDP. Hence, harnessing these benefits entails performing data preparation within the SSD, despite the SSD’s limited hardware resources. Unfortunately, this is ch allenging because performing random accesses to large amounts of data inside the SSD is inefficient due to costly conflicts, e.g., in NAND flash chips and channels. Although modern SSDs have an internal DRAM buffer, its capacity is small [^6] A modern SSD typically has 1-GB DRAM for each TB capacity. For example, a 4-GB LPDDR4 DRAM is used in a 4-TB SSD (samsung860pro, )., mostly filled with mapping metadata, and its bandwidth is constrained by its single-channel interface (zou2022assasin, ). Across our datasets, we observe that the state-of-the-art genomic decompression (i) requires random accesses to large amounts of data (up to 26 GB), significantly larger than the internal DRAM, and (ii) as shown by our motivational analysis, even on our high-end system with eight DRAM channels, its performance saturates after a specific thread count due to main memory bandwidth bottlenecks.
Our goal in this work is to improve the performance and energy efficiency of genome analysis by mitigating the data preparation bottleneck. We aim for the proposed design to achieve high performance and energy efficiency, high compression ratios comparable to state-of-the-art genomic compressors, and to be lightweight for seamless integration with broad ranges of genome analysis systems.